OpenStudy (anonymous):

HELP ME REALLY NEED HELP FREE MEDAL AND FAN IF YOU HELP ME Daniel is looking over some data regarding the temperatures of countries and how the temperature relates to the number of mosquitoes. The data are represented by the exponential function f(x) = 2x, where x is the number of mosquitoes. Explain how he can convert this equation to a logarithmic function when the temperature is 64 degrees.

OpenStudy (anonymous):

f(x)=2^x

OpenStudy (anonymous):

could u explain

OpenStudy (anonymous):

it's asking you to convert the equation that they give you to a logarithmic function when he temperature is 64 and explain it

OpenStudy (anonymous):

oh kk thnks

OpenStudy (anonymous):

@ganeshie8 ,@Hero ,@Preetha

OpenStudy (anonymous):

help me please

OpenStudy (jewlzme17):

Do you know how to do logarithms?

OpenStudy (anonymous):

not really but if you explain it to me ill get it

OpenStudy (anonymous):

im not the best at logarithms

OpenStudy (anonymous):

Is anyone really good at logarithm

OpenStudy (jewlzme17):

okay so logarithms is and equation put into log form. To put an equation into log form you go in a circle of the equation and that is where you place everything. For example y=b^x is equivalent to log_b(y)=x so F(x)=2^x would be log_2f(x)=x

OpenStudy (jewlzme17):

That is regardless of the temp. Does f(x) represent the temp? That is what is confusing me. If so then just input 64 into the f(x) position.

OpenStudy (anonymous):

yes f(x) represents the temp

OpenStudy (jewlzme17):

okay so then just input that. So it would come out to log_264=x. Mind you the the underbar isn't there. That is just to show that the 2 is small.

OpenStudy (jewlzme17):



OpenStudy (anonymous):

ok am i suppose to find the x

OpenStudy (jewlzme17):

No that is the logarithmic equation right there. Not unless your teacher wants you to solve it but that is that logarithmic equation!

OpenStudy (anonymous):

ok thank you so much I get it

OpenStudy (jewlzme17):

No problem! :)